OpenStudy (anonymous):

Compute the standard deviation for the set of sample data 90, 81, 73, 73, 81

OpenStudy (anonymous):

doing this by hand is annoying

OpenStudy (anonymous):

to do so you first have to find the average of the set

OpenStudy (amistre64):

and recall that sample data has to /(n-1)

OpenStudy (amistre64):

something to do with the degree of freedom

OpenStudy (anonymous):

the average is 79.6

OpenStudy (amistre64):

2(81) + 2(73) + 90 ---------------- = mean 4

OpenStudy (anonymous):

now take the square of each number subtracted from the mean

OpenStudy (anonymous):

u lose a degree of freedom coz u estimate population mean by taking the mean of the sample.

OpenStudy (amistre64):

2(81+73) + 90 2(154)+ 90= 398 99.5 is the mean if i did it right ----- 4|398 36 38 36 20

OpenStudy (anonymous):

so (90-79.6)^2+(81-79.6)^2+...

OpenStudy (anonymous):

amistre64, the mean is 79.6, she is correct

OpenStudy (anonymous):

it wouldnt make sense for the mean to be more than any of the numbers in the set...

OpenStudy (amistre64):

i see that :) means i typoed or mismathed lol

OpenStudy (amistre64):

im thinking of the n-1 in a different spot .. arent i

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

perhaps? \[\sqrt{\frac{\sum{(x-\bar x)^2}}{n-1}}\]

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

so i got 207.14

OpenStudy (anonymous):

you should have gotten 199.2

OpenStudy (amistre64):

my ti83 pops out Sx = 7.057 to check with

OpenStudy (anonymous):

ya, thats what i got in excel too

OpenStudy (anonymous):

from the 199.2, you divide by n-1 where n is the number of numbers in the series

OpenStudy (anonymous):

then square root

OpenStudy (zarkon):

I would use the formula \[s=\frac{n\sum x^2-(\sum x)^2}{n(n-1)}\]

OpenStudy (zarkon):

it's quicker

OpenStudy (amistre64):

i still wonder why; since sd is the average distance from the mean, why we dont take the mean of the absolute values of x - mean ... i fhtat makes sense

OpenStudy (zarkon):

square root of that

OpenStudy (anonymous):

its actually longer to do by hand...

OpenStudy (zarkon):

\[s=\sqrt{\frac{n\sum x^2-(\sum x)^2}{n(n-1)}}\]

OpenStudy (zarkon):

quicker than using the definition by hand

OpenStudy (anonymous):

not for me, guess im weird

OpenStudy (zarkon):

you are then ;)